Solve using elimination, thank you!
x - 0.5y = -3
2x - y = 6

Answer by Alan3354(69443)

Solve using elimination, thank you!
x - 0.5y = -3
2x - y = 6
=========================
2x - y = -6 ---- 1st eqn times 2
2x - y = 6
------------------ Subtract
0 = -12
--> inconsistent, no solution.
If you graph them you get parallel lines.
No intersection, no solution.